40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 255 of 329  Not logged in ELOG logo
ID Date Author Type Category Subject
  3777   Mon Oct 25 11:39:06 2010 JenneUpdateSUSMagnets glued to PRM
This morning I glued the magnets to the PRM. Now we wait, and tomorrow afternoon (at the earliest), Suresh and I can balance the PRM.
Attachment 1: StatusTable.png
StatusTable.png
  3776   Mon Oct 25 02:25:21 2010 KojiUpdatePSLQuarter Wave Plate Measurements

Q1. Suppose the laser beam has a certain (i.e. arbitrary) polarization state but contains only TEM00. Also suppose the PSB is perfect (reflect all S and transmit all P). What results do you expect from your expereiment?

Q2. Suppose the above condition but the PBS is not perfect (i.e. reflects most of S but also small leakage of P to the reflection port.) How are the expected results modified?

Q3. In reality, the laser may also contain some thing dirty (e.g. deporarization in the laser Xtal, higher order modes in a certain polarization but different from the TEM00's one, etc). What actually is the cause of 170mW rejection from the PBS? Can we improve the transmitted power through the PBS?

Q4. Why is the visibility for the lambda/4 with 330deg better than the one with 326deg? Yes, as I already explained to Kevin, I suppose it was caused by the lack of the data points in the wider angle ranges.

Quote:

I measured the reflected power from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations.

The optimum angles that minimize the reflected power are 330° for the quarter wave plate and 268° for the half wave plate.

The following data was taken with 2.102 A laser current and 32.25° C crystal temperature.

For each of five quarter wave plate settings around the optimum value, I measured the reflected power from the PBS with an Ophir power meter. I measured the power as a function of half wave plate angle five times for each angle and averaged these values to calculate the mean and uncertainty for each of these angles. The Ophir started to drift when trying to measure relatively large amounts of power. (With approximately 1W reflected from the PBS, the power reading rapidly increased by several hundred mW.) So I could only take data near the minimum reflection accurately.

The data was fit to P = P0 + P1*sin^2(2pi/180*(t-t0)) with the angle t measured in degrees with the following results:

lambda/4 angle (°) t0 (°) P0 (mW) P1 (mW) chi^2/ndf V
318 261.56 ± 0.02 224.9 ± 0.5 2016 ± 5 0.98 0.900 ± 0.001
326 266.07 ± 0.01 178.5 ± 0.4 1998 ± 5 16.00 0.918 ± 0.001
330 268.00 ± 0.01 168.2 ± 0.3 2119 ± 5 1.33 0.926 ± 0.001
334 270.07 ± 0.02 174.5 ± 0.4 2083 ± 5 1.53 0.923 ± 0.001
342 273.49 ± 0.02 226.8 ± 0.5 1966 ± 5 1.41 0.897 ± 0.001

where V is the visibility V = 1- P_max/P_min. These fits are shown in attachment 1. We would like to understand better why we can only reduce the reflected light to ~150 mW. Ideally, we would have V = 1. I will redo these measurements with a different power meter that can measure up to 2 W and take data over a full period of the reflected power.

 

  3775   Mon Oct 25 02:23:47 2010 KojiConfigurationelogELOG 2.8.0
When I push the reply button, the raw html shows up in the edit window and have to use HTML to write the entry.
Does this happen only to me???


Quote:
I stopped the ELOG and restarted us on 2.8.0.

To make sure nothing got lost, I killed the old process, copied over the logbooks/, themes/, and elogd.cfg to the new 2.8.0/ directory before starting the new Daemon.

I encountered the same Administrator bug as Joe had before. I delete all the old Admin passwords to bypass the issue.

To restart the ELOGD on NODUS, you now type '/cvs/cds/caltech/elog/start-elog.csh'.
I also added ELOG to the man pages in /usr/local/man/ on nodus by putting the *.1 files in man1/ and the *.8 files into man8/.
  3774   Mon Oct 25 02:14:38 2010 KojiSummaryLockingsetup for green beat

- What is the actual photocurrent for the beam1 and beam2? We don't care how much power do you have before the BS, but care how much current do you have on the PD.

- How much is the visibility? There is mismatching of the beams. i.e. The beam diameter looked quite different. Also the beams are not TEM00 but have fringes probably comes from the TT mirrors. You maybe able to measure the visibility by the DC output, making the beat freq go through df=0 slowly.

- What is the measured gain of the RF amp? Does it include the voltage division by the output/input impedance?

---------------------

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

  3773   Sun Oct 24 19:55:50 2010 kiwamuUpdateElectronicslonely RF amplifier on ITMX table
(Rana, Kiwamu)

Last Friday we found a lonely RF amplifier ZHL-3A on the ITMX table.
When we found him we were very sad because he's been setup unacceptably


For example, the signal input was disconnected although a 24V DC was still applied. So he has been making just a heat for a long time.
The power connector was a BNC style which is not official way.
The leg of a decoupling capacitor attached to the DC connector was apparently broken and etc,..

We salvaged him and then cleaned up those cables and the DC power supply.

We don't say like 'don't make a temporary setup', but please clean up them after finishing the work every time.
  3772   Sun Oct 24 19:23:41 2010 ranaConfigurationelogELOG 2.8.0
I stopped the ELOG and restarted us on 2.8.0.

To make sure nothing got lost, I killed the old process, copied over the logbooks/, themes/, and elogd.cfg to the new 2.8.0/ directory before starting the new Daemon.

I encountered the same Administrator bug as Joe had before. I delete all the old Admin passwords to bypass the issue.

To restart the ELOGD on NODUS, you now type '/cvs/cds/caltech/elog/start-elog.csh'.
I also added ELOG to the man pages in /usr/local/man/ on nodus by putting the *.1 files in man1/ and the *.8 files into man8/.
  3771   Sun Oct 24 18:06:35 2010 kiwamuSummaryLockingsetup for green beat

green_beat_note.png

(notes on signal level)

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

 

The expectation was calculated by using the following simple math.

                       SIGNAL = A x Z x G_RF x sqrt(P1 / 2) x sqrt (P2 / 2)

 where A is the responsibility of the PD and Z is the trans impedance gain. G_RF is a gain of the RF amplifier.

The laser powers of green beams, P1 and P2, are divided by 2 due to a beam splitter. 

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

 

  3770   Sat Oct 23 14:42:01 2010 yutaSummaryCDSdamped MC suspensions

After replacing filters, MC suspensions damped  last night.

Further measurement next time.

Attachment 1: dam.png
dam.png
  3769   Sat Oct 23 03:36:05 2010 yutaUpdateCDSfixed filters for C1SUS, C1RMS, C1MCS

(Joe, Yuta)

Summary:

 This Monday, MC suspension damping got something wrong.
 We started to check filters and found that digital filters were wrong because of mis-conversion from old filter files to new files.
 We converted the file again, and with mutual understanding between Foton and us, we finally got correct filters(I hope!).

What we did:
 1. Merged filter files in old /cvs/cds/caltech/chans/ directory into C1SUS.txt(BS,ITMX,ITMY), C1RMS.txt(PRM,SRM), C1MCS.txt(MC123).

 2. Rebuilt the RT models in order to get a correct filter file header(they have list of filter modules).

 3. Concatenate the header with filter design part which we got from step1.

 4. Replaced 'N 2048' with 'N 16384'
   It replaces sampling rate of "XXSEN"s.

Basically, step1-4 was the same with what we did last time. We didn't changed the fitler coefficients, so Foton somehow changed the original filter design.
So, this time, we

 5. Deleted coefficients like we did on Tuesday (see elog #3774).

But Foton couldn't read the file correctly this time. Foton seemed to be unbeatable.
Even if we replaced the sampling rate, Foton kept saying 2048! (This is maybe because Foton's default value is 2048Hz. Everytime Foton notice some editting in the file, he destroys everything. He hates editting)
The problems were always associated with the sampling rate, so

 6. Got back to step4 and undo-ed the replacement.

 7. Foton could read it this time, so I changed the sampling rate one by one using Foton GUI.

 8. Checked filters using Foton's Bode Plot.(Not for all, but some that had problem before)

 9. Splitted SDSEN filters to SDSEN, SUSSIDE, and SDCOIL.

 10. Put some missing whitening filters, and 28HzELPs.
   BS, PRM, SRM didn't have any 28HzELP for SDCOIL.
   ITMX and ITMY SDCOIL had SimDW/InvDW which doesn't make sense(SIDEs don't have analog DW). So, I deleted and replaced with 28HzELP.

F2A issue:
 We failed in sending F2A filters to new filter files.
 These are a little bit complicated because TO_COIL_X_X filters were named ULPOS,URPOS etc before.
 Also, MC3 didn't have any F2As, so maybe we should but the same F2As as MC1/2.
 Note that every F2As are different, and TO_COIL matrix have UL,UR,LL,LR order(not same as INMATRIX).
 Also, SRM had f2pv instead of F2A!

Next work:
 - Check whole filters by actually measuring transfer function between SENs and COILs.
 - Damp MC suspentions, and lock MC.
 - Measure openloop TF and compare with the designed.


How do you read a Foton filter file:
 When you open up a Foton filter file, you see filters like this.

################################################################################
### modulename                                                               ###
################################################################################
# SAMPLING modulename samplingrate
# DESIGN   modulename n filterdesign
# DESIGN   modulename n filterdesign
###                                                                          ###
modulename   n xy z      v      w filtername                        gain    a1     a2    b1     b2
modulename   n xy z      v      w filtername                        gain    a1     a2    b1     b2
                                                                            a1     a2    b1     b2
n: filter number
 0 for FM1, 1 for FM2, ... , 9 for FM10

x: Input Switching setting
 1 Always On
 2 Zero History

y: Output Switching setting
 1 Immediately
 2 Ramp
 3 Input Crossing
 4 Zero Crossing

z: number of filters cascaded.

v: if y=2, (Ramp Time(sec))*(samplingrate)
   if y=3 or 4, Tolerance

w: (Timeout(sec))*(samplingrate)

 Note that v and w are changed when sampling rate is changed.

 Transfer function will be;
  H(1/z)=G*(1+b1/z+b2/z/z)/(1+a1/z+a2/z/z)
  z=exp(s/fs)

 where fs is the sampling frequency.

Reference:
 Kiwamu Izumi: "Notes about Digital Filters," http://tamago.mtk.nao.ac.jp/izumi/green/DigitalFilter.pdf

  3768   Sat Oct 23 02:25:49 2010 KevinUpdatePSLQuarter Wave Plate Measurements

Quote:

[Koji and Kevin]

We measured the reflection from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations. Before the measurement we reduced the laser current to 1 A, locked the PMC, and recorded 1.1 V transmitted through the PMC. During the measurements, the beam was blocked after the faraday isolator. After the measurements, we again locked the PMC and recorded 1.2 V transmitted. The current is now 2.1 A and both the PMC and reference cavities are locked.

I will post the details of the measurement tomorrow.

I measured the reflected power from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations.

The optimum angles that minimize the reflected power are 330° for the quarter wave plate and 268° for the half wave plate.

The following data was taken with 2.102 A laser current and 32.25° C crystal temperature.

For each of five quarter wave plate settings around the optimum value, I measured the reflected power from the PBS with an Ophir power meter. I measured the power as a function of half wave plate angle five times for each angle and averaged these values to calculate the mean and uncertainty for each of these angles. The Ophir started to drift when trying to measure relatively large amounts of power. (With approximately 1W reflected from the PBS, the power reading rapidly increased by several hundred mW.) So I could only take data near the minimum reflection accurately.

The data was fit to P = P0 + P1*sin^2(2pi/180*(t-t0)) with the angle t measured in degrees with the following results:

lambda/4 angle (°) t0 (°) P0 (mW) P1 (mW) chi^2/ndf V
318 261.56 ± 0.02 224.9 ± 0.5 2016 ± 5 0.98 0.900 ± 0.001
326 266.07 ± 0.01 178.5 ± 0.4 1998 ± 5 16.00 0.918 ± 0.001
330 268.00 ± 0.01 168.2 ± 0.3 2119 ± 5 1.33 0.926 ± 0.001
334 270.07 ± 0.02 174.5 ± 0.4 2083 ± 5 1.53 0.923 ± 0.001
342 273.49 ± 0.02 226.8 ± 0.5 1966 ± 5 1.41 0.897 ± 0.001

where V is the visibility V = 1- P_max/P_min. These fits are shown in attachment 1. We would like to understand better why we can only reduce the reflected light to ~150 mW. Ideally, we would have V = 1. I will redo these measurements with a different power meter that can measure up to 2 W and take data over a full period of the reflected power.

Attachment 1: fits.png
fits.png
  3767   Fri Oct 22 23:29:48 2010 yutaUpdateCDSfixed output filter switching

(Joe, Yuta)

Summary:
 This evening, we fixed output filter analog/digital switchings in Simulink logic, but they didn't actually switched filters.
 That was because what we thought BO0 was BO2 and BO2 was BO0.
 We re-wired, re-labeled the cables.
 We checked it by using a LED(see "test configuration" in this wiki page).
 Also, we checked the Simulink connection by probing MAX333A in SOS Dewhitening Board located at 1X4-8-11B, which switches SIDE coils for all optics.

Next work:

 - Currently, FM10 in UL/UR/LR/LL/SDCOIL filters switch analog/digital, but let's make FM9 do the switching.(See schematic in elog #3758)
 - Check all the logic by measuring transfer functions(maybe single frequency measurement is enough. we can use tdsdmd etc to automate).

  3766   Fri Oct 22 21:28:27 2010 AidanUpdateLockingGreen era: found a green beat note finally

Nice work!

Quote:

finally we found it !

 green_beatnote.jpg

 

  3765   Fri Oct 22 19:53:27 2010 yutaSummaryCDSconversion failure in digital filters

(Rana, Joe, Yuta)

We now understand that we never succeeded in converting old fiter files to new filter files.

For example, we just changed the sampling rate with coefficients remained and foton confused, or we forgot to rename some of the module names(ULSEN -> SUS_MC1_ULSEN) ......

This is why we sometimes damped and sometimes didn't, depending on filter switches. New filter files has been always wrong.

So, we started to convert them again.
We have to figure out how to convert the files that Foton accepts correctly.

  3764   Fri Oct 22 18:22:27 2010 AlbertoUpdateElectronicsEpic Takeover

Quote:

As the suspension work winds down (we'll be completely done once the ETMs arrive, are suspended, and then are placed in the chambers), I'm going to start working on the RF system. 

Step 1: Figure out what Alberto has been up to the last few months.

Step 2: Figure out what still needs doing.

Step 3: Complete all the items listed out in step 2.

Step 4: Make sure it all works.

Right now I'm just starting steps 1 & 2.  I've made myself a handy-dandy wiki checklist: RF Checklist.  Hopefully all of the bits and pieces that need doing will be put here, and then I can start checking them off. Suggestions and additions to the list are welcome.

 There's also a page dedicated to the progress in the PD upgrade process:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/RF_System/Upgraded_RF_Photodiodes

There you can find a pdf document with my notes on that.

  3763   Fri Oct 22 18:15:21 2010 kiwamuUpdateLockingGreen era: found a green beat note finally

finally we found it !

 green_beatnote.jpg

  3762   Fri Oct 22 16:59:21 2010 JenneUpdateElectronicsEpic Takeover

As the suspension work winds down (we'll be completely done once the ETMs arrive, are suspended, and then are placed in the chambers), I'm going to start working on the RF system. 

Step 1: Figure out what Alberto has been up to the last few months.

Step 2: Figure out what still needs doing.

Step 3: Complete all the items listed out in step 2.

Step 4: Make sure it all works.

Right now I'm just starting steps 1 & 2.  I've made myself a handy-dandy wiki checklist: RF Checklist.  Hopefully all of the bits and pieces that need doing will be put here, and then I can start checking them off. Suggestions and additions to the list are welcome.

  3761   Fri Oct 22 15:06:43 2010 JenneUpdateSUSArts and Crafts!

This afternoon I epoxied the guiderod and wire standoff to the new PRM.  I also epoxied the magnets that Suresh picked out to the dumbbell standoffs.  We'll let them all cure over the weekend, and then I'll glue the magnets to the optic on ~Monday.

Notes about the epoxy: 

Previously, we had been using the "AN-1" epoxy, which is gray, with a clear hardener.  Bob recommended we switch to "30-2", which is clear with clear, and has been chosen for use in aLIGO.  Both were vacuum approved, but the 30-2 has gone through ~2 months of testing at the OTF (Optics Test Facility?) over in Downs under vacuum, to check the level of outgassing (or really, non-outgassing).

The 30-2 is less viscous than the AN-1, and it takes less glue to do the same job, so we should keep that in mind when applying the epoxy.  When I put the glue next to the guiderod and standoff, it got wicked along the length of each rod, which is good.  I can't reach the whole length of the rod with my glue applicator because the fixture holding them in place blocks access, so the wicking is pretty handy.

 

I've also added the updated version of my Status Table for the suspensions.

Attachment 1: StatusTable.png
StatusTable.png
  3760   Fri Oct 22 03:37:56 2010 KevinUpdatePSLQuarter Wave Plate Measurements

[Koji and Kevin]

We measured the reflection from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations. Before the measurement we reduced the laser current to 1 A, locked the PMC, and recorded 1.1 V transmitted through the PMC. During the measurements, the beam was blocked after the faraday isolator. After the measurements, we again locked the PMC and recorded 1.2 V transmitted. The current is now 2.1 A and both the PMC and reference cavities are locked.

I will post the details of the measurement tomorrow.

  3759   Fri Oct 22 01:23:13 2010 KojiUpdatePSLFound the beat at 1064nm

[Koji / Suresh]

We worked on the 1064 beating a bit more.

- First of all, FSS and FSS SLOW servo were disabled not to have variating Xtal temp for the PSL.

- The PSL Xtal temp (T_PSL) was scanned from 22deg-45deg while we search the Xtal temp (T_Xend) for the Xend laser to have the beat freq well low (f<30MHz).
The pumping current for each laser was I_PSL = 2.101 [A] and I_Xend = 2.000 [A]

For a certain T_PSL, we found multiple T_Xend because the freq of the laser is not a monotonic function of the Xtal temperature. (see the innolight manual).

T_Xend to give us the beating was categorized in the three sets as shown in the figure. The set on "curve2" is the steadiest one. (Use this!)
The trends were quite linear but the slope was slightly off from the unity.

- T_PSL was scanned to see the trend of the PMC output.

The PMC was sometimes locked to the mode with lower transmission (V_PMCT ~ 3.0V).
When T_PSL ~ 31deg we consistently locked the PMC at higer transmission (V_PMCT ~ 5.3V).

At the moment we decided the operating point of T_PSL = 32.25 deg, V_PMCT = 5.34, where we found the beat at T_Xend=38.28deg.

- We cleaned up the PSL table more than how it was. Returned the tools to their original places.
The X-end laser was shut down and was left on the PSL table.

NEXT:
Kiwamu can move the X-end laser to the Xend and realign it.
Then we should be able to see the green beating quite easily.

Attachment 1: 101021_beat.pdf
101021_beat.pdf
  3758   Fri Oct 22 00:59:01 2010 yutaUpdateCDSnew MEDM screens, new filter banks(whitening and dewhitening related)

(Joe, Yuta)

Summary:
 No more discrimination for SIDE!
 We had all SIDE filters in SDSEN, so we splitted into SDSEN, SUSSIDE, and SDCOIL just like other DOFs. (Not SIDECOIL. SDCOIL from now on!)
 Also, there was a confusion in output filters, so we fixed the filter bank(dewhitening and 28HzELP).
 More than that, I found that "13HzCheby" in SUSPOS, SUSPIT, SUSYAW was actually doing ~1.6Hz Chebyshev, so I fixed it. (No wonder we had no damping when turing "13HzCheby" on!)
 Our new MEDM screen is Attachment #1.

Input filter design:
 Every optics, every OSEM has same analog whitening filter.
 So;

digital analog
30,100:3 30,100:3
ON OFF
OFF ON

Output filter design:
 For every SIDE coils and MC1, MC3 coils, they have analog 28Hz elliptic LPFs.
 So;
digital analog
28HzELP 28HzELP
ON OFF
OFF ON

 For MC2(and other main optics) UL/UR/LR/LL coils, they have analog dewhitening filters.
 So;
digital analog
InvDW SimDW DW
ON ON OFF
ON OFF ON

 InvDW and SimDW were in FM10 and FM9, but I swapped them to make it more consistent.
 Now, FM10 switches analog/digital output filiters for every optics.

  Let's put 28HzELP in FM9 so that FM9 switches analog/digital for every optics.

Quote:

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Next work:
 - Fix Simulink connection and logic so that analog/digital switching go correctly.
 - Check if the switching are correct for all channels by measuring transfer functions.
 - Currently, we are focusing on MCs, but we have to do the same fixing for other optics, too.
 - More than setting the right filter TF, there are switching settings like Zero History, Ramp, Zero Crossing...... We should investigate that.
 - Actually, TF of analog DW in D000183 doesn't look like the same with digital SimDW. Why??

Attachment 1: HappySideMEDM.png
HappySideMEDM.png
  3757   Thu Oct 21 21:35:16 2010 yutaUpdateCDSfixed whitening filter switches

(Joe, Yuta)

Summary:
 We found some mistakes in whitening filter switches yesterday(see elog #3748).
 One was the mis-connection between LL/LR and the second was the bit invert.
 So, we fixed it and checked that they are now switching correct by probing MAX333A.

Why we made mistake:
1. LL/LR mis-connection
  Simply, Simulink model was wrong.
  It occurred because UL/UR/LR/LL order is currently arbitrary and confusing.
  For now, we just swap the connection(Attached #1).
  In the future, we should fix all orders corresponding to the actual wiring order(UL > LL > UR > LR).

2. Bit invert
  It occurred from counter-intuitive output of Contec 32 BO(See Attachment #2(taken from this wiki page)).
  If "1," current goes from A to B, so MAX333A input will be 0V, which means analog WF is on.
  If "0," no current, so MAX333A input will be +15V, which means analog WF is off(bypassed).

Next work:
 - There are no digital WF in SDSEN inputs. So, we have to put them.
 - Dewhitening filters are totally messed up now. Switches are wrong, some optics have digital DW but some doesn't, some have 28Hz elliptic LPF.......
     We have to grand unify them.

Attachment 1: LLLR.png
LLLR.png
Attachment 2: Contec23BO.png
Contec23BO.png
  3756   Thu Oct 21 19:10:39 2010 AidanUpdatePSLFound the beat at 1064nm

Quote:

[Koji Suresh]

We found the beat at 1064nm. T(PSL)=26.59deg, T(X-end)=31.15deg.

The X-end laser was moved to the PSL table.

The beating setup was quickly constructed with mode matching based on beam profile measurements by the IR cards.
We used the 1GHz BW PD, Newfocus #1611-FS-AC.

As soon as we swept the Xtal temp of the X-end laser, we found the strong beating.

 

  3755   Thu Oct 21 18:45:50 2010 KojiUpdatePSLFound the beat at 1064nm

[Koji Suresh]

We found the beat at 1064nm. T(PSL)=26.59deg, T(X-end)=31.15deg.

The X-end laser was moved to the PSL table.

The beating setup was quickly constructed with mode matching based on beam profile measurements by the IR cards.
We used the 1GHz BW PD, Newfocus #1611-FS-AC.

As soon as we swept the Xtal temp of the X-end laser, we found the strong beating.

  3754   Thu Oct 21 15:59:28 2010 kiwamuUpdateLockingno green beat note : details

Last night I tried searching for a beat note signal with two different PD trans impedance gains.

Although I didn't find a  beat note signal.

 


 

- (1. trans impedance gain = 2400)

 I started with a trans impedance resistance of R=2.4k, which is 10 times bigger resistance than the original.

The total PD gain should be about 960 [V/W] theoretically if we assume the responsibility of the PD is 0.4 [A/W].

Then I checked the bandwidth of the RFPD using Jenne laser.

The bandwidth was about 30MHz, which is 3 times narrower than the original. And it agrees with our expectation.

As Koji and I mentioned at the last weekly meeting, the cut off frequency of an RFPD follows inverse square root of the trans impedance resistance R.

 

 bw.png

where C is a capacitance of the photo diode. (See this)

I was expecting the signal level of  -50 dBm / rtHz with a 23dB RF amplifier, assuming the line width of the signal is 10kHz.

 

 


- (2. trans impedance gain = 240) 

 I also tried it with the original trans impedance gain (see this entry). 

 R = 240 [Ohm]

 G = 96 [V/W]

 BW = 100 [MHz]  (I didn't measure it in this time)

 expected signal level = -70 dBm/rtHz


Quote:

 Still I didn't see any beat note signals..

 

  3753   Thu Oct 21 13:05:19 2010 KojiSummaryComputerselog change and rossa tex

This is cool though the projector is flashing the blue screen alternately.

I gave the dual head video card (ATI RADEON something) to Yuta a month ago.
It is on the top of Zita. This would make the things more fun.

Quote:

Yuta has also successfully set up zita to run the projector, so we should clear our some of the boxes and bookshelves in that area so that the projection can be larger.

 

  3752   Thu Oct 21 12:15:02 2010 ranaUpdatePEM6.9 Mag EQ in Gulf of California
Magnitude 6.9
Date-Time
Location 24.843°N, 109.171°W
Depth 10 km (6.2 miles) set by location program
Region GULF OF CALIFORNIA
Distances 105 km (65 miles) S of Los Mochis, Sinaloa, Mexico
125 km (75 miles) SW of Guamuchil, Sinaloa, Mexico
140 km (85 miles) NE of La Paz, Baja California Sur, Mexico
1200 km (740 miles) WNW of MEXICO CITY, D.F., Mexico
Location Uncertainty horizontal +/- 6.1 km (3.8 miles); depth fixed by location program
Parameters NST=187, Nph=187, Dmin=843.1 km, Rmss=1.17 sec, Gp=133°,
M-type=teleseismic moment magnitude (Mw), Version=6
Source
  • USGS NEIC (WDCS-D)
Event ID us2010crbl
  3751   Thu Oct 21 10:44:56 2010 steveMetaphysicsTreasure1987 supernova tapes plus....?

I'm cleaning out to make room for our new optical cabinet. Are we keeping these? There are  ~20  pieces of 10" od 1" wide tapes and large number of cassettes.

AJW,  Zucker,  Stuart A and Koji were notified in this matter.

Alan suggested to save data of Bruce Allen paper of observation of binary neutron stars in the 40m on 1994 November 14-20 and save back up tapes of his period in the 40m.

Mike: reels are not readable any more, it is time to let go

Attachment 1: P1060932.JPG
P1060932.JPG
  3750   Thu Oct 21 05:51:10 2010 kiwamuUpdateLockingno green beat note

 Still I didn't see any beat note signals..

 With a help from Suresh, Yuta and Rana, I tried searching for the green beat note by changing the temperature of the X end NPRO.

The noise level after it goes through an RF amplifier (G=23dB) was about -70dBm at 50MHz.

This noise may cover the beat note signal.

I am going to post some details later.

  3749   Thu Oct 21 01:11:48 2010 ranaSummaryComputerselog change and rossa tex

I deleted a bunch of categories from the elogd.cfg file. Not every kind of locking and activity gets its own activity. All of the things related to length sensing and control should be LSC. If there's a high pollen count its under PEM. etc., etc., etc.

I decided that the 'tetex' distribution which comes with CentOS is total BS and I removed it from rossa using 'yum erase tetex'. I installed TexLive using the instructions on the web; its much better, actually works, and also has RevTex 4.1 which allows Jenne and I to write our paper.

Eventually, we'll standardize our workstation setups.

Yuta has also successfully set up zita to run the projector, so we should clear our some of the boxes and bookshelves in that area so that the projection can be larger.

  3748   Wed Oct 20 21:43:25 2010 yutaUpdateCDSchecking whitening filters for MCs and messed up

(Joe, Yuta)

Background:
 We found that the damping servo for MC suspensions somehow worked when we turned off the 13Hz Chebyshev filters.
 But that does not meet our satisfaction, so we started checking every components.
 First of all is the whitening filters.
 If we turn on a digital whitening filter(WF), corresponding analog WF should turn off, and vice versa.

Reference:
 See DCC #D000210 for the analog circuit of WF. WF has 3Hz zero, 30Hz pole, 100Hz pole. MAX333A bypasses analog WF when supplied +15V(HIGH).

What we did:
 1. Compared the transfer function between MC2_SUSPOS_EXC and MC2_ULSEN_IN1 when digital WF on and off. When digital is on/off, analog should be off/on, so there should be difference but couldn't see.

 2. We went through the simulink model and found 2 mistakes in the logic. One is the conflict with other optics. Even if we turn on/off digital WF of MC2, it didn't switched analog WF of MC2. Two is the additional bit invert (but it turned out to be our misunderstanding).

 3. We (thought we) fixed it, rebuild it, and measured the TF again.(Attachment #1)

[Attached #1]
 The red/blue line is when digital WF is on/off. Blue should be bigger(+10dB @ 10Hz according to (analog) WTF) than red, but it was the opposite.

 4. To confirm that they are doing the opposite, I checked MAX333A input(pin#1,10,11,20) in "SUS PD Whitening Board" at 1X5-1-8B (which is for MC3) and found that switching is opposite. When I turned off/on the digital WF, MAX333A input was +15V/0V. It should be 0V/+15V.

 5. Also, I found that LLSEN digital WF switch switches LRSEN analog WF and vice versa.

[Attached #2]
 Transfer functions between MC2_SUSPOS_EXC and MC2_LRSEN_IN1 with;
   Red: LR digital off, LL digital on
   Blue: LR digital off, LL digital off
   Green: LR digital on, LL digital off
 As you can see, LL switch is the one which switches LR analog WF now.

Conclusion:
 Currently, digital WF on/off corresponds to analog WF on/off.
 Also, LL/LR digital WF switch is LR/LL analog WF switch now.


Next work:
 - fix the simulink models (or wiring)
 - check dewhitening filters

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Attachment 1: MC2ULSEN.png
MC2ULSEN.png
Attachment 2: MC2LRSEN.png
MC2LRSEN.png
  3747   Wed Oct 20 21:33:11 2010 KevinUpdatePSLQuarter Wave Plate Optimization

[Suresh and Kevin]

We placed the quarter wave plate in front of the 2W laser and moved the half wave plate forward. To make both wave plates fit, we had to rotate one of the clamps for the laser. We optimized the angles of both wave plates so that the power in the reflection from the PBS was minimized and the transmitted power through the faraday isolator was maximized. This was done with 2.1 A injection current and 38°C crystal temperature.

Next, I will make plots of the reflected power as a function of half wave plate angle for a few different quarter wave plate rotations.

  3746   Wed Oct 20 18:17:35 2010 Suresh, JenneUpdateSUSPRM assembly

We have positioned the guide rod and the wire-stand-off on the optic in the axial direction. 

We have selected six magnets whose magnetic strength is +/-5% of their mean strength (180 Gauss).  The measurement was made as follows:

1) each magnet was placed on its  end, on the top of a beaker held upside down. 

2) The Hall probe was placed directly under the magnet touching the glass from the other side (the inside of the beaker). 

This ensures that the relative position of the magnet and the probe remains fixed during a measurement.  And ensures that their separation is the same for each of the magnets tested. 

With this procedure the variation in the measured B field is less than +/- 10% in the sample of magnets tested.

  3745   Wed Oct 20 16:41:41 2010 steveOmnistructureGeneral laser chiller leaving for MIT
Attachment 1: P1060929.JPG
P1060929.JPG
  3744   Wed Oct 20 01:22:18 2010 yutaUpdateCDSfixed filters for MCs, but no damping

(Rana, Yuta)

Summary:

 The damping servo for MCs has not been working since Monday.
 We already found that some filter coefficients were wrong.(see elog #3742)
 So, we fixed it (just for MCs now).
 But still doesn't work.

What we did:
Fixed the filters for MC suspensions;
 1. In /cvs/cds/rtcds/caltech/c1/chans/ directory, we replaced C1MCS.txt with ./filter_archive/c1mcs/C1MCS_101019_101927.txt.
  This is the archive of the filter bank for MC suspensions, when filter designs were correct(when we hadn't opened it with foton yet).

 2. Deleted all the filter coefficients.
   By using emacs magical C-<space> C-x r k.

 3. Loaded with foton to get correct coefficients.

 4. Reloaded coefficients for C1MCS.

We now have the correct filters, but the damping still doesn't work.

 5. To confirm that the filters are now correct and the model is working correctly, we measured the transfer function between ULSEN input and ULCOIL output and compared with the calculated TF from the filter bank file (Attachment #1) .

 6. To calculate the designed function, we used the following matlab scripts;
    /cvs/cds/caltech/users/rana/mat/utilities/FotonFilter.m        # takes the foton file and returns the TF
    /cvs/cds/caltech/users/rana/mat/utilities/readFilterFile.m    # reads the foton file
    /cvs/cds/caltech/users/yuta/scripts/sentocoil.m                   # calculates the total TF from SEN to COIL

Result:
 As you can see from the attachment, filters seem fine now. (The red line is the measured and the blue line is the calculated function in the plot)
 But the damping (for POS, PIT, YAW) still doesn't work. Why????

Next work:

 - see if coils are pushing correctly
 - push cable connectors further!
 - check input and output filters (whitening, dewhitening)
 - when fixed, use my fully updated QAdjuster.py for adjusting Q for multiple channels automatically!!

Attachment 1: Screenshot.png
Screenshot.png
  3743   Tue Oct 19 22:37:28 2010 KojiUpdatePSLPSL table cleaning up

I cleaned up the scattered tools, optics, and mounts of the PSL table. I gathered those stuffs at the two coners.

At the end of the work I scanned the table with an IR viewer. (This is mandatory)
I put some beam block plates to kill weak stray beams.

One thing I like to call the attention is:

I found that some beam blocks were missing at around the PBSs just after the laser source.
Those PBSs tend to reject quite a lot of beam power
--- no matter how the HWPs/QWPs are arranged.

--- even at the backward side.
(remember that we have a faraday there.)

Particularly, there was no beam block at the forward rejection side of the first PBS where we dump the high power beam.

Be careful. 

  3742   Tue Oct 19 21:10:27 2010 yutaUpdateCDSbad filter coefficients for every suspensions!

(Koji, Yuta)

Summary:
 The damping servos of the MC suspensions has not been working since yesterday. They had worked on Friday.
 We found that some of the filter coefficients somehow changed from the correct value during dealing with the sampling frequency shift (from 2048Hz to 16384Hz). (see elog #3659)

What we did:
We thought that the problem occurred from the CDS update on Monday. So, we restored them using svn.
 1. Went to /opt/rtcds/caltech/c1/core/advLigoRTS and ran svn update using the revision number 2125.
   svn update -r 2125
 2. Rebuilt all the front ends.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/core/advLigoRTS
   make clean-SYSNAME
   make SYSNAME
   make install-SYSNAME 
   (where SYSNAME=c1x02,c1sus,c1mcs,c1rms)
 3. Restarted the front end models.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/scripts
   sudo startc1SYS 
   (where SYS=sus,mcs,rms)
 4. Restarted mx_streams.
   sudo /etc/restart_streams
See this wiki page for the restart procedures.

At this point we judged that the realtime system and "dataviewer" worked fine (by poking some suspensions / by measuring PSDs/TFs of the signals).

However, just restoring the RT system didn't fix the damping servos.

After that, we measured the transfer functions of the servo filters using DTT(diaggui on rosalba).
 5. The filter at MC1_SUSPOS, "3:0.0" should have a cut-off frequency at 3Hz, but it was around 0.3Hz.
 6. We used foton in order to look at the filter bank files (in /cvs/cds/rtcds/caltech/c1/chans). Then we found that the filter design was somehow changed to
   zpk([0],[0.375],2.66666,"n")
  instead of the correct one
   zpk([0],[3],0.333333,"n")

Why the filter designs changed?:
 We suspect that the filter designs were changed because of the replacement of sampling frequency in filter bank files(see elog #3659). We only replaced the lines like
  # SAMPLING SUS_MC3_SUSPOS 2048
 to
  # SAMPLING SUS_MC3_SUSPOS 16384
 and didn't changed the coefficients.
 This may caused a problem when opening the files with foton, and foton changed 3Hz to 0.375Hz (2048/16384) and other coefficients.

Note:
 The damping servo for SIDE has been working for every MCs. POS, PIT, YAW are the ones not working.

Next work:
 - Fix filters for every suspensions.
   There are backup files in /cvs/cds/rtcds/caltech/c1/chans/filter_archive directory, so this should help.
   But I don't think just fixing filters would fix the damping servo because the filter design of MC suspensions were changed this morning and the damping had worked until Friday.

  3741   Tue Oct 19 15:14:51 2010 JenneUpdateSUSPRM (little) update

[Jenne, Suresh]

We've aligned the guiderod and wire standoff to the PRM, each partly.  They have both been aligned to the correct distance above the scribe lines, but they have not yet been centered forward/backward along the thickness of the optic.  So, we're working on it...

  3740   Tue Oct 19 00:24:07 2010 DmassOmnistructureElectronicsMassive restocking of the 40m

I had a number of delinquent items on the sign out list from the 40m. I returned about half, and ordered replacements for most of the other half.

I put the photodiodes on the SP table, and the 560 on the electronics  bench.

  3739   Mon Oct 18 22:11:32 2010 Joonho LeeSummaryElectronicsCCD cables for input signal

Today I checked all the CCD cables which is connected input channels of the VIDEOMUX.

Among total 25 cables for output, 12 cables are type of RG59, 4 cables are type of RG58, and 9 cables are of unknown type.

The reason I am checking the cables is for replacing the cables with impedance of 50 or 52 ohm by those with impedance of 75 ohm.

After I figures out which cable has not proper impedance, I will make new cables and substitute them in order to match the impedance, which would lead to better VIDEO signal.

 

Today, I check the cables in similar way as I did the last time.

I labeled all cables connected to input channels of VIDEO MUX and disconnect all of them since last time it was hard to check every cable because of cables too entangled.

Then I checked the types of all cables and existing label which might designate where each cable is connected to.

After I finished the check, I reconnected all cables into the input channel which each of cable was connected to before I disconnected.

 

4 cables out of 25 are type of RG58 so expected to be replace with cable of type RG59.

9 cables out of 25 are of unknown type. These nine cables are all orange-colored thick cables which do not have any label about the cable characteristic on the surface.

The result of observation is as follows.

Note that type 'TBD-1' is used for the orange colored cables because all of them look like the same type of cable.

 

Channel number where its signal is coming type
1 C1:IO-VIDEO 1 MC2 TBD-1
2 FI CAMERA 59
3 PSL OUTPUT CAMERA 59
4 BS  C:1O-VIDEO 4 TBD-1
5 MC1&3 C:1O-VIDEO 5 59
6 ITMX C:1O-VIDEO 6 TBD-1
7 C1:IO-VIDEO 7 ITMY TBD-1
8 C1:IO-VIDEO 8 ETMX TBD-1
9 C1:IO-VIDEO 9 ETMY TBD-1
10 No cable is connected
(spare channel)
 
11 C1:IO-VIDEO 11 RCR 59
12 C1:IO-VIDEO RCT 59
13 MCR VIDEO 59
14 C1:IO-VIDEO 14 PMCT 59
15 VIDEO 15 PSL IOO(OR IOC) 59
16 C1:IO-VIDEO 16 IMCT TBD-1
17 PSL CAMERA 58
18 C1:IO-VIDEO 18 IMCR 59
19 C1:IO-VIDEO 19 SPS 59
20 C1:IO-VIDEO 20 BSPO TBD-1
21 C1:IO-VIDEO 21 ITMXPO TBD-1
22 C1:IO-VIDEO 22 APS1 59
23 ETMX-T 58
24 ETMY-T 58
25 POY CCD VIDEO CH25 58
26 OMC-V 59

Today I could not figure out what impedance the TBD-1 type(unknown type) has.

Next time, I will check out the orange-colored cables' impedance directly and find where the unknown output signal is sent. Any suggestion would be really appreciated.

  3738   Mon Oct 18 18:33:46 2010 KojiSummaryCOCSummary of the main mirrors & their phasemap measurement

I have made a summary web page for the 40m upgrade optics.

https://nodus.ligo.caltech.edu:30889/40m_phasemap/

I made a bunch of RoC calculations along with the phase maps we measured.
Those are also accommodated under this directory structure.

Probably.... I should have used the wiki and copy/paste the resultant HTML?

  3737   Mon Oct 18 18:00:36 2010 KojiUpdateSUSOld PRM, SRM stored, new PRM drag wiped

- Steve is working on the storage shelf for those optics.

- PRMU002 was chosen as it has the best RoC among the three.

Quote:

[Jenne, Suresh]

We've put the old PRM and SRM (which were living in a foil house on the cleanroom optical table) into Steve's nifty storage containers.  Also, we removed the SRM which was suspended, and stored it in a nifty container.  All 3 of these optics are currently sitting on one of the cleanroom optical tables.  This is fine for temporary storage, but we will need to find another place for them to live permanently.  The etched names of the 3 optics are facing out, so that you can read them without picking them up.  I forgot to note the serial numbers of the optics we've got stored, but the old optics are labeled XRM ###, whereas the new optics are labeled XRMU ###. 

Koji chose for us PRMU 002, out of the set which we recently received from ATF, to be the new PRM.  Suresh and I drag wiped both sides with Acetone and Iso, and it is currently sitting on one of the rings, in the foil house on the cleanroom optical table.

We are now ready to begin the guiderod gluing process (later tonight or tomorrow).

 

  3736   Mon Oct 18 17:16:30 2010 JenneUpdateSUSOld PRM, SRM stored, new PRM drag wiped

[Jenne, Suresh]

We've put the old PRM and SRM (which were living in a foil house on the cleanroom optical table) into Steve's nifty storage containers.  Also, we removed the SRM which was suspended, and stored it in a nifty container.  All 3 of these optics are currently sitting on one of the cleanroom optical tables.  This is fine for temporary storage, but we will need to find another place for them to live permanently.  The etched names of the 3 optics are facing out, so that you can read them without picking them up.  I forgot to note the serial numbers of the optics we've got stored, but the old optics are labeled XRM ###, whereas the new optics are labeled XRMU ###. 

Koji chose for us PRMU 002, out of the set which we recently received from ATF, to be the new PRM.  Suresh and I drag wiped both sides with Acetone and Iso, and it is currently sitting on one of the rings, in the foil house on the cleanroom optical table.

We are now ready to begin the guiderod gluing process (later tonight or tomorrow).

  3735   Mon Oct 18 15:33:00 2010 josephb, alexUpdateCDSPossibly broken timing cards

It looks as though we may have two IO chassis with bad timing cards.

Symptoms are as follows:

We can get our front end models writing data and timestamps out on the RFM network.

However, they get rejected on the receiving end because the timestamps don't match up with the receiving front end's timestamp.  Once started, the system is consistently off by the same amount. Stopping the front end module on c1ioo and restarting it, generated a new consistent offset.  Say off by 29,000 cycles in the first case and on restart we might be 11,000 cycles off.  Essentially, on start up, the IOP isn't using the 1PPS signal to determine when to start counting.

We tried swapping the spare IO chassis (intended for the LSC) in ....

# Joe will finish this in 3 days.

# Basically, in conclusion, in a word, we found that c1ioo IO chassis is wrong.

  3734   Mon Oct 18 11:22:13 2010 JenneUpdateComputersShame on people not elogging! FrameFiles backups not working.

On the one hand, SHAME ON ALL PEOPLE WHO DON'T ELOG THINGS, such as the moving of scripts directories (it was a pain to figure out that that's part of why the backup scripts are broken).  On the other hand, the moving of the scripts directories brought to light a critical problem in the backup scheme. None of the frame files have been backed up since Joe replaced fb40m with fb, on ~23 Sept (I think).

What went down:

The frame builder was replaced, and no backup script was started up on the new machine.  Sadface.  Crontab doesn't work yet on the new machine, and also the 'ssh-add' commands give an error:

controls@fb /cvs/cds/rtcds/caltech/c1/scripts/backup $ ssh-add ~/.ssh/id_rsa
No such file or directory
controls@fb /cvs/cds/rtcds/caltech/c1/scripts/backup $ ssh-add ~/.ssh/backup2PB 
No such file or directory

Thus, I know that the backup was never running on the new fb.  However, the check-er script runs on nodus, and looks at the logfile, and since there was no script running, it wasn't adding to the log file, so the last log was an "Okay, everything worked" entry.  So, the check-er script kept sending me daily emails saying that everything was okie-dokey. 

Since all of the scripts were moved (Joe said this happened on Friday, although there's no elog about it), the check-er script, and all of the rest of the backup scripts point to the wrong places (the old scripts/backup directory), so I didn't receive any emails about the backup either way (usually it at least sends a "Hey, I'm broken" email).  This clued me in that we need to check things out, and I discovered that it's all gone to hell.

Since I can't add the ssh clients to the new fb, I can't actually log in to the backup computers over in Powell-Booth to check when the last legitimately successful backup was. But I suspect it was just before the fb was replaced.

So, we need to get Crontab up and running on the new Frame Builder machine so that we can run cron jobs, and we also need to figure out this backup hullabaloo.  I think I'll email / call Dan Kozak over in downs, who was talking about upgrading our backup scheme anyway.

 

  3733   Mon Oct 18 09:01:48 2010 steveHowToGeneralHow not to

This BNC cable is crying for help. Please do not do this to me.  It should be reported  to the  abused center.Throw this cable into the garbage now.

Attachment 1: P1060926.JPG
P1060926.JPG
  3732   Mon Oct 18 08:44:36 2010 ranaUpdateCDSslow DAQ channels added

Why EXCMON? I think that we should modify the filter coefficients so that the decimation filter that makes OUT16 is a 2nd order Butterworth at 1 Hz instead of whatever bogus thing it is now. Then we can delete the EXCMON from the DAQ and add either OUTPUT or OUTMON. That way we will have a low frequency signal (OUT16) and a sort of aliased RMS (OUTMON).

  3731   Fri Oct 15 22:16:22 2010 kiwamuUpdateSUSwhat's wrong with MC2

[Yuta, Suresh, Rana and Kiwamu]

 The DC alignment problem of MC2 was fixed.

There were some loosely connected cables on the backside of a VME rack which contains the MC2 SOS driver.

We pushed those connectors to make them tightly connected.  And then the problem disappeared.

 


(voltage unbalance on coils)

 Before fixing it Yuta opened the satellite box and measured the voltage across the coils using a voltmeter.

At that time UL and LR showed 20 times smaller voltages than that of the other two when we moved the DC alignment slider from min. to max. on the medm screen.

This behavior is exactly consistent with the wired motion of a beam spot which we saw when we were aligning MC2. 

 

(diagnostic using optical lever)

 After pushing the connectors, we made an optical lever using a red laser pointer in order to check the actual motion of MC2.

We confirmed that MC2 respond correctly to the alignment slider.

Quote:

 2010:Oct.14th:

It turned out that the DC alignment of MC2 from epics doesn't helathily work.

For example, the pitch slider does drive the yaw alignment as well.

 

  3730   Fri Oct 15 21:25:23 2010 SureshUpdateIOO2W NPRO laser output power drop question

  The power meter used in the measurements of elog entries 2822, 3698 and 3709 was the Ophir PD300-3W.  This power head has several damaged patches  and a slight movement of the laser spot changes the reading considerably.  To verify I checked the power out with another power meter (the Vector S310) and found that there is no significant variation of the power output with the temperature of the laser.  And the power at 2.1A of diode current is 2W with 10% fluctuation arising from slight repositioning of the laser head.  There are regions of the Ophir PD300 which show the laser power to be about 1.9W.

  3728   Fri Oct 15 15:32:35 2010 josephbUpdateCDSslow DAQ channels added

I added all the _OUT16, _INMON, _EXCMON channels associated with the BS, ITMX, and ITMY channels in SUS_SLOW.ini in the /opt/rtcds/caltech/c1/chans/daq directory.  Similarly, I added the channels associated with MC1, MC2, and MC3 to MCS_SLOW.ini and those associated with PRM and the SRM to RMS_SLOW.ini. 

Lines pointing to these files were then added to the /opt/rtcds/caltech/c1/target/fb/master file and the frame builder restarted.  It took about 4 tries before the frame builder stayed up using ./daqd -c ./daqdrc.

I generated the .ini files with a python script.  Its at /opt/rtcds/caltech/c1/scripts/daqscripts/create_inmon_out16_daq_ini.py. It checks the C0EDCU.ini to find what slow channels already exist, then goes through a medm directory given at the command line looking for all _OUT16, _INMON, and _EXCMON channels, adding them if they aren't already accounted for, and then writing out the new file to the location of your choice.

  3727   Fri Oct 15 06:31:52 2010 kiwamuUpdateGreen LockingPSL beat note setup: part II

 I made some KAIZENs (what does kaizen mean ? ) for the PSL green setup.

I replaced the lenses for the modematching of the two green lights at the PSL table, and the beams now look pretty identical.

Also I tuned the temperature setpoint of the doubling crystal and eventually the green light increased to 14 uW at the PSL table.

Once I finish the modification of the RFPD tomorrow, I am going to search for the beat note signal.

 


( details )

 - In-vac green mirrors

   I found one of the green steering mirror, which stands at the corner of the MC table, was clipping the green light.

  So I steered another mirror, which sends the beam to the clipping mirror after the downward periscope.

 I touched also the last steering mirror in the OMC chamber to correct the alignment.

 

 - temperature of the doubling crystal

   I took a quick temperature scan in order to find an optimum point for the crystal temperature.

  The scan was performed by just turning the heater off after I heated up the crystal up to 40 deg. 

  Using the NewPort power meter I found the optimum point around 37.3 deg. So I set the temperature to that point.

 

 - mode matching lenses

 As written in this entry , Kevin and I had put some lenses to make the two green beam almost the same size at the RFPD.

 But today while I was checking these mode-profiles by using a sensor card, I found they were not so matched.

 Therefore I replaced these lenses to match them more.

The idea of this replacement is t to let them have a long Rayleigh range, such that they can efficiently and easily interfere because of the flatness of their wave fronts along a distant.

 For the green light from the chamber, I put one more lens to form a Keplerian beam shrinker (see here about the Keplerian lens configuration).

 They look pretty identical now.

 

ELOG V3.1.3-